OpenStudy (anonymous):

which expression is equivalent to 2/3(4x + 9) 8/3x +6 8/3x + 9 2/3x + 6 2/3x + 9

OpenStudy (anonymous):

please help ASAP

OpenStudy (anonymous):

1st one. Wanna learn why?

OpenStudy (anonymous):

(2/3) times 4x=(8/3x). 2/3 times 9=6. therefore the answer is 8/3x + 6

OpenStudy (anonymous):

please help ASAP

OpenStudy (anonymous):

please help ASAP

OpenStudy (anonymous):

please help ASAP

OpenStudy (anonymous):

please help ASAP

OpenStudy (anonymous):

I literally just gave you the answer lol